The length of a rectangular box is 4 times its width, and the height is 6ft more than its width. The volume of the box is 52 ft cubed. I can't figure it out because you need a graphing calculator that graphs it and than you click zero to get the answer. That where I'm puzzled. And round two decimal places
Answer by rfer(16322)

4x*x*x+6=52
4x^3=46
x^3=11.5
x=2.26 width
4x=9.04 length
x+6=8.26 height